# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v1

@@ 4.1.1 Margin Impact and Results
    The Dunn Index improved significantly with higher margins, especially for failure SETs, which achieved their highest separability at a margin of 1.5. This indicated that increasing the margin value results in more well-defined clusters, with greater distances between the success and failure groups.
@@
In terms of epoch loss, we did not observe significant changes with varying margin values Fig.8. The overall epoch loss progression remained relatively stable, suggesting that while higher margins did not heavily impact convergence or loss reduction during training, they did have a positive effect on the quality of clustering in the latent space. In the cluster analysis below, we examined how these higher margin values enhanced episode separability, supporting generalisation and task-based decision-making. Figure 8 illustrates the training loss for different margin values.
Figure 8: Training loss for various margin values. Higher margins improve clustering quality without significantly affecting convergence.
@@ 4.1.2 Visualization of Outcome Embeddings
Figure 9(a) shows the K-means clustering of SET embeddings, color-coded by task. The results highlight SETLE’s ability to capture task-level distinctions and result-specific information.
(a) Task-level clusters: K-means Clustering of SET Encodings (PCA Reduced). Each point represents a SET in the CREATE environment, color-coded by task.
@@ 4.1.3 Loss Functions Impact and Results
To assess the effectiveness of the hybrid loss in improving representation compared to the standard triplet loss, we visualised the SET embeddings generated by SETLE under both loss functions. Each point represents a SET, colour-coded by task, with successful episodes marked as $\square$ and unsuccessful episodes marked as $\times$. The visualisations are provided in Figures 10(a) and 10(b), where subfigure (a) shows the results with triplet loss and subfigure (b) shows the results with hybrid loss.
@@
The hybrid loss significantly enhanced the representation of inter-task relationships while maintaining strong intra-task and outcome-level differentiation, by using cross-view contrastive learning in addition to triplet-based alignment.
@@
4.1.4 Comparison on CREATE environment
@@
4.2 S2: Ablation Studies Emphasising the Model’s Design
@@ 1. Random State Removal
Design: Randomly removed a fixed proportion (e.g., 10%, 20%, 40%) of the states in the graph. Encode the modified SET using SETLE and compare clustering quality. Below we presented the results of removing 40% of the states, having observed that for lower values the results were insignificant.
@@ 2. Sequentiality Disruption
Remarks: Sequential disruption reduced the ability to encode cumulative effects, leading to embeddings were are less meaningful and cohesive.
@@ 3. Flattened Representations
Remarks: Flattening the graph significantly degraded the quality of the learnt task representations, highlighting the importance of relational data for meaningful task dynamics (see Table 3).
Results:

### v2

@@ 4.3.1 Margin Impact and Results
    We observed an appreciable improvement in the Dunn Index with higher margins, especially for failure SETs, which achieved their highest separability at a margin of 1.5. This indicated that increasing the margin value results in more well-defined clusters, with greater distances between the success and failure groups.
@@
In terms of epoch loss, we did not observe significant changes (see Section 4.6) with varying margin values Fig.7. The overall epoch loss progression remained relatively stable, suggesting that while higher margins did not heavily impact convergence or loss reduction during training, they did have a positive effect on the quality of clustering in the latent space. In the cluster analysis below, we examined how these higher margin values enhanced episode separability, supporting generalisation and task-based decision-making. It is important to note the trade-off between clustering quality and training stability. While our statistical analysis (Section 4.6) confirmed that lower margin values lead to a significantly lower final training loss (t(5)= -2.83, p=0.037), the results in Table 1 clearly show that higher margin values are superior for achieving the primary goal of this encoder: creating a well-separated latent space. This demonstrates a classic hyperparameter trade-off where a stronger constraint (a higher margin) improves the desired structural outcome at the cost of a slightly more difficult optimization process.
Figure 7: Comparison of training loss curves for high-margin ($\alpha$=1.5, red dashed line) and low-margin ($\alpha$=0.1, blue solid line) configurations of the loss. Both models demonstrate stable convergence, indicating that while the margin value is critical for the final separability of embeddings (see Table 1), it does not significantly hinder the model’s ability to learn (Section 4.6).
@@ 4.3.2 Visualization of Outcome Embeddings
Figure 8(a) shows the K-means clustering of SET embeddings, colour-coded by task. The results highlight SETLE’s ability to capture task-level distinctions and result-specific information.
(a) Task-level clusters: K-means Clustering of SET Encodings (PCA Reduced). Each point represents a SET in the CREATE environment, colour-coded by task.
@@ 4.3.3 Loss Functions Impact and Results
To assess the effectiveness of the hybrid loss in improving representation compared to the standard triplet loss, we visualised the SET embeddings generated by SETLE under both loss functions. Each point represents a SET, colour-coded by task, with successful episodes marked as $\square$ and unsuccessful episodes marked as $\times$. The visualisations are provided in Figures 9(a) and 9(b), where subfigure (a) shows the results with triplet loss and subfigure (b) shows the results with hybrid loss.
@@
The hybrid loss considerably enhanced the representation of inter-task relationships while maintaining strong intra-task and outcome-level differentiation, by using cross-view contrastive learning in addition to triplet-based alignment.
@@
4.4 Comparison on CREATE environment
@@
4.4.1 S2: Ablation Studies Emphasising the Model’s Design
@@ 1. Random State Removal
Design: Randomly removed a fixed proportion (e.g., 10%, 20%, 40%) of the states in the graph. Encode the modified SET using SETLE and compare clustering quality. Below we presented the results of removing 40% of the states, having observed that for lower values the results were not statistically reliable.
@@ 2. Sequentiality Disruption
Remarks: Sequential disruption reduced the ability to encode cumulative effects, leading to embeddings that are less meaningful and cohesive.
@@ 3. Flattened Representations
Remarks: Flattening the graph drastically degraded the quality of the learnt task representations, highlighting the importance of relational data for meaningful task dynamics (see Table 3).